# Leave-One-Out Analysis for Nonconvex Robust Matrix Completion with General Thresholding Functions

### v1

@@ 1 Introduction
To the best of our knowledge, there is no leave-one-out analysis for nonconvex methods targeting the RMC problem. In a closely related work [9], theoretical recovery guarantee is improved for a convex program by building a connection between the optimal solution of the convex program with the iterates of a nonconvex algorithm. The nonconvex algorithm can be viewed as a projection-free version of the method used in [29], except that the sorting operation is replaced by the soft-thresholding operation. However, the nonconvex algorithm in [9] is not a practical one, since it requires initialization at the ground truth and is only used to verify certain optimal conditions. That being said, it motivates us to derive the theoretical guarantee of certain nonconvex methods for RMC based on the leave-one-out analysis technique, thus achieving a projection and sample splitting free result. Overall, our main contributions can be summarized as follows.
  • We consider a nonconvex method which alternates between SVP for the low rank estimation and entrywise thresholding for the sparse outlier estimation. The theoretical recovery guarantee of the nonconvex method has been established based on the leave-one-out analysis technique for a class of thresholding functions, including for example soft-thresholding and SCAD. To the best of our knowledge, this is the first leave-one-out analysis for nonconvex methods targeting the RMC problem.
  • Though we have adopted an analysis route that is similar to that used in [11], the effect of the sparse outliers has to be carefully examined. It also requires us to extend the proofs for positive-definite matrices to general matrices and provide bounds for new terms which appear due to asymmetry. Moreover, our result is also applicable for low rank matrix completion, in which case we have managed to reduce the sample complexity of SVP from $O(\frac{\kappa^{6}\mu^{4}r^{6}\log n}{n})$ in [11] to $O(\frac{\kappa^{6}\mu^{3}r^{3}\log n}{n})$ using different technical tools. Technical differences between proofs in our paper and that in [11] have been briefly discussed at the beginnings of Sections 4 and 5.
@@ 1.1 Assumptions and Notations
Assumption 1.
Assume $L^{\star}\in\mathbb{R}^{n\times n}$ is $\mu$-incoherent, i.e.,
@@
Assumption 2.
Each entry is independently observed with probability $p$, i.e.,
@@
Assumption 3.
Denote by $\Omega_{S^{\star}}$ the support of the outliers in $\Omega$. Assume that each observed entry is independently corrupted by an outlier with probability $\alpha$, i.e.,
$\mathbb{P}\left\{(i,j)\in\Omega_{S^{\star}}~{}|~{}(i,j)\in\Omega\right\}=\alpha.$
Remark 1. Assumptions 1 and 2 are standard assumptions in the matrix completion literature. Assumption 1, together with Assumption 3, can be considered as separation conditions on $(L^{\star},S^{\star})$ [5, 9]. Combining Assumptions 2 and 3, one can show that, with high probability [^1] Throughout the paper, this terminology means that the event occurs with probability at least $1-C_{1}n^{-C_{2}}$ for some constants $C_{1},C_{2}>0$ and $C_{2}$ sufficiently large., there are at most $2\alpha p$-fraction of outliers per row and column, i.e., $\mathcal{P}_{\Omega}\left(S^{\star}\right)$ is $2\alpha p$-sparse.
@@ 2.1 Algorithm and Main Result
Theorem 1.
$p\geq\frac{C_{\emph{sample}}}{\gamma^{2}}\cdot\frac{\kappa^{6}\mu^{3}r^{3}\log n}{n},\quad\alpha\leq\frac{c_{\emph{outlier}}}{\kappa^{3}\mu^{2}r^{2}}\cdot\frac{\gamma}{C_{\emph{thresh}}}$
and
for iteration $0\leq t\leq T$, where $T=n^{O(1)}$.
Remark 3. Note that our result is also applicable for the noiseless case, i.e. the low rank matrix completion problem. In this case, we have improved the sample complexity of SVP from $O(\frac{\kappa^{6}\mu^{4}r^{6}\log n}{n})$ in [11] to $O(\frac{\kappa^{6}\mu^{3}r^{3}\log n}{n})$.
@@ 3 Proof Outline
$\max_{0\leq i\leq 2n}\left\|E^{t-1,i}\right\|_{2}:=\left\|E^{t-1,\infty}\right\|_{2}$ $\leq\frac{1}{C_{0}}\cdot\left(\frac{1}{\kappa^{2}\mu r}\sigma_{r}^{\star}\right)\gamma^{t}$ (3a)
$\max_{0\leq i\leq 2n}\left\|\Delta^{t,i}\right\|_{2,\infty}:=\left\|\Delta^{t,\infty}\right\|_{2,\infty}$ $\leq\frac{5C_{10}}{C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t}$ (3b)
@@
$\leq$ $\frac{5C_{10}}{C_{0}}\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t}\cdot\left(2+\frac{5C_{10}}{C_{0}}\right)\sqrt{\frac{\mu r}{n}}\cdot\left(1+\frac{1}{C_{0}}\right)\sigma_{1}^{\star}+\frac{\mu r}{n}\cdot\frac{7}{C_{0}}\left(\frac{1}{\kappa\mu r}\sigma_{r}^{\star}\right)\gamma^{t}$
$\leq$ $\frac{12C_{10}}{C_{0}}\cdot\frac{1}{\kappa\mu r}\left(\frac{\mu r}{n}\sigma_{1}^{\star}\right)\gamma^{t}\leq\left(\frac{\mu r}{n}\sigma_{1}^{\star}\right)\gamma^{t},$
@@ 4 Initialization and Base Case
  • We use Lemma 4 instead of [11, Lemma 21] to bound $\left\|\mathcal{H}_{\Omega}\left(L^{\star}\right)\right\|_{2}$, and use Bernstein’s inequality applied to vectors instead of scalars to bound $\left\|e_{m}^{T}\mathcal{H}_{\Omega}^{(-i)}\left(L^{\star}\right)V^{1,m}\right\|_{2}$, both yielding better bounds.
@@
Step 1: The Bound For $\left\|E^{0,\infty}\right\|_{2}$. First consider $\left\|E_{1}^{0,i}\right\|_{2}$. Recall from Remark 1 that with high probability, $\mathcal{P}_{\Omega}\left(S^{\star}\right)$ is $2\alpha p$-sparse. Conditioning on this event, one has
$\left\|E_{1}^{0,i}\right\|_{2}=$ $\left\|p^{-1}\mathcal{P}_{\Omega}^{(-i)}\left(S^{0,i}-S^{\star}\right)\right\|_{2}\leq p^{-1}\cdot\left(2\alpha pn\right)\cdot C_{\text{thresh}}\frac{\mu r}{n}\sigma_{1}^{\star}\leq\frac{\gamma}{2C_{0}}\cdot\frac{1}{\kappa^{2}\mu r}\sigma_{r}^{\star},$ (4)
@@
$\alpha\leq\frac{1}{4C_{0}}\frac{1}{\kappa^{3}\mu^{2}r^{2}}\cdot\frac{\gamma}{C_{\text{thresh}}}.$
@@
$\leq$ $2c_{4}\sqrt{\frac{\mu r\log n}{np}}\sigma_{1}^{\star}\leq\frac{\gamma}{2C_{0}}\cdot\frac{1}{\kappa^{2}\mu r}\sigma_{r}^{\star},$
@@
$p\geq\frac{16c_{4}^{2}C_{0}^{2}}{\gamma^{2}}\cdot\frac{\kappa^{6}\mu^{3}r^{3}\log n}{n}.$
@@
$\left\|E^{0,i}\right\|_{2}\leq\left\|E_{1}^{0,i}\right\|_{2}+\left\|E_{2}^{0,i}\right\|_{2}\leq\frac{1}{C_{0}}\cdot\left(\frac{1}{\kappa^{2}\mu r}\sigma_{r}^{\star}\right)\gamma.$ (6)
@@
$\left\|R\right\|_{2}\leq\left\|\Sigma^{\star}H^{1,i}-G^{1,i}\Sigma^{\star}\right\|_{2}\cdot\left\|\left(\Sigma^{\star}\right)^{-1}\right\|_{2}\leq\frac{4}{C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\right)\gamma.$
@@
$\left\|T_{1}\right\|_{2}\leq\frac{4}{C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma.$
@@
$\left\|\left(\Sigma^{1,i}\right)^{-1}-\left(\Sigma^{\star}\right)^{-1}\right\|_{2}=\max_{1\leq k\leq r}\left|\frac{1}{\sigma_{k}^{1,i}}-\frac{1}{\sigma_{k}^{\star}}\right|=\max_{1\leq k\leq r}\frac{\left|\sigma_{k}^{1,i}-\sigma_{k}^{\star}\right|}{\sigma_{k}^{1,i}\sigma_{k}^{\star}}\leq\frac{2}{C_{0}}\cdot\left(\frac{1}{\kappa^{2}\mu r}\right)\gamma\cdot\frac{1}{\sigma_{r}^{\star}}.$
@@
$\left\|T_{2}\right\|_{2}\leq\frac{2}{C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma.$
@@
$\left\|T_{3}\right\|_{2}\leq\frac{20}{C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma.$
@@
$\leq$ $\frac{1}{2C_{0}}\frac{\gamma}{\kappa^{2}\mu r}\sigma_{r}^{\star}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{1,\infty}\right\|_{2,\infty}\right),$
@@
$\leq$ $\frac{C_{10}}{2c_{4}C_{0}}\frac{\gamma}{\kappa^{2}\mu r}\sigma_{r}^{\star}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{1,\infty}\right\|_{2,\infty}\right),$
@@
$\left\|T_{4}\right\|_{2}\leq\frac{\gamma}{C_{0}}\left[\left(1+\frac{C_{10}}{c_{4}}\right)\left(\frac{1}{\kappa^{2}\mu r}\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{1,\infty}\right\|_{2,\infty}\right)+\left\|D^{1,\infty}\right\|_{\mathrm{F}}\right].$
@@
$\left\|\Delta^{1,\infty}\right\|_{2,\infty}\leq\frac{17+\frac{C_{10}}{c_{4}}}{C_{0}}\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma+\frac{1+\frac{C_{10}}{c_{4}}}{C_{0}}\left\|\Delta^{1,\infty}\right\|_{2,\infty}+\frac{1}{C_{0}}\left\|D^{1,\infty}\right\|_{\mathrm{F}}.$ (8)
@@
$\leq$ $\frac{1}{2C_{0}}\frac{\gamma}{\kappa^{2}\mu r}\sigma_{r}^{\star}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{1,\infty}\right\|_{2,\infty}\right),$
@@
$\leq$ $\frac{1}{2C_{0}}\frac{\gamma}{\kappa^{2}\mu r}\sigma_{r}^{\star}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{1,\infty}\right\|_{2,\infty}\right),$
@@
$B_{2}=\left\|e_{m}^{T}\mathcal{H}_{\Omega}\left(L^{\star}\right)V^{1,m}\right\|_{2}\leq\frac{C_{10}}{2c_{4}C_{0}}\frac{\gamma}{\kappa^{2}\mu r}\sigma_{r}^{\star}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{1,\infty}\right\|_{2,\infty}\right),$
@@
$\leq$ $\frac{1}{2C_{0}}\frac{\gamma}{\kappa^{2}\mu r}\sigma_{r}^{\star}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{1,\infty}\right\|_{2,\infty}\right),$
@@
$\left\|\Delta^{1,\infty}\right\|_{2,\infty}$ $\leq\frac{17+\frac{C_{10}}{c_{4}}}{C_{0}}\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma+\frac{1+\frac{C_{10}}{c_{4}}}{C_{0}}\left\|\Delta^{1,\infty}\right\|_{2,\infty}$
@@
$\left(1-\frac{2+2C_{10}}{C_{0}}\right)\left\|\Delta^{1,\infty}\right\|_{2,\infty}\leq$ $\frac{18+2C_{10}}{C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma,$
$\left\|\Delta^{1,\infty}\right\|_{2,\infty}\leq$ $\frac{18+2C_{10}}{C_{0}-2-2C_{10}}\cdot\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma.$
@@
$\left\|\Delta^{1,\infty}\right\|_{2,\infty}\leq\frac{5C_{10}}{C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma.$

### v2

@@ 1 Introduction
1.1 Main Contributions
To the best of our knowledge, there is no leave-one-out analysis for nonconvex methods targeting the RMC problem. In a closely related work [9], theoretical recovery guarantee is improved for a convex program by building a connection between the optimal solution of the convex program with the iterates of a nonconvex algorithm. The nonconvex algorithm can be viewed as a projection-free version of the method used in [29], except that the sorting operation is replaced by the soft-thresholding operation. However, the nonconvex algorithm in [9] is not a practical one, since it requires initialization at the ground truth and is only used to verify certain optimal conditions. That being said, it motivates us to derive the theoretical guarantee of certain nonconvex methods for RMC based on the leave-one-out analysis technique, thus achieving a projection and sample splitting free result. Overall, the novelty of this paper are summarized as follows.
  1. 1.
    We consider a nonconvex method which alternates between SVP for the low rank estimation and entrywise thresholding for the sparse outlier estimation. The theoretical recovery guarantee of the nonconvex method has been established based on the leave-one-out analysis technique for a class of thresholding functions, including for example soft-thresholding and SCAD. It is also worth pointing out that our analysis can handle any sparse enough outliers, without requiring the support and signs of the outliers to be random, as assumed in [9]. To the best of our knowledge, this is the first leave-one-out analysis for nonconvex methods targeting the RMC problem.
  2. 2.
    When applying our result to the matrix completion problem, our sample complexity improves substantially upon the one in [11] over the dependency on $\kappa$, $\mu$ and $r$, i.e., from $O(\frac{\kappa^{6}\mu^{4}r^{6}\log n}{n})$ to $O(\frac{\kappa^{4}\mu^{3}r^{3}\log n}{n})$.
For the analysis, as inspired by [11], we rewrite the update of the algorithm as a perturbation around the ground truth and then establish the contraction of $\|E^{t-1,\infty}\|_{2}$, $\|\Delta^{t,\infty}\|_{2,\infty}$, and $\|D^{t,\infty}\|_{F}$ (see Section 3 for the definitions) by induction and the leave-one-out analysis. That being said, we would like to emphasize that the proof details are quite technically involved and new techniques have to be developed to achieve the improved analysis and to handle the outliers.
  • Firstly, to achieve the improved sample complexity, instead of first bounding $\|\Delta^{t,\infty}\|_{2,\infty}$ and then bounding $\|D^{t,\infty}\|_{F}$ based on $\|\Delta^{t,\infty}\|_{2,\infty}$, we have established two inequalities involving $\|\Delta^{t,\infty}\|_{2,\infty}$ and $\|D^{t,\infty}\|_{F}$ simultaneously. This allows us to relax the requirement on the distance to the ground truth when performing the analysis. In addition, the extension of the proofs in [11] for positive-definite matrices to general matrices needs to be done properly, including providing bounds for new terms appeared due to asymmetry.
  • Secondly, while the outlier terms in the initialization can be controlled easily by the sparsity of the outliers, the outlier terms in the induction steps need to be bounded carefully. In particular, our analysis does not require the pattern and the signs of the outliers to be random (as is assumed in [9]), but is applicable for any sort of outliers.
  • Last but not least, more effective technical tools have been used in the initialization and induction steps (for example using [7, Lemma 2] instead of [11, Lemma 21]) and new lemmas have also been developed (for example Lemmas 11 and 13).
@@ 1.2 Assumptions and Notations
Assumption 1. Assume $L^{\star}\in\mathbb{R}^{n\times n}$ is $\mu$-incoherent, i.e.,
@@
Assumption 2. Each entry is independently observed with probability $p$, i.e.,
@@
Assumption 3. Denote by $\Omega_{S^{\star}}$ the support of the outliers in $\Omega$. It is assumed that $\mathcal{P}_{\Omega}\left(S^{\star}\right)$ is $2\alpha p$-sparse, i.e., it has no more than $2\alpha pn$ nonzero entries per row and column.
Remark 1. Assumptions 1 and 2 are standard assumptions in the matrix completion literature. Assumption 1, together with Assumption 3, can be considered as separation conditions on $(L^{\star},S^{\star})$ [5, 9]. One can see that if each observed entry independently has probability $\alpha$ to be corrupted by an outlier, Assumption 3 can be satisfied with high probability [^1] Throughout the paper, this terminology means that the event occurs with probability at least $1-C_{1}n^{-C_{2}}$ for some constants $C_{1},C_{2}>0$ and $C_{2}$ sufficiently large.. However, Assumption 3 can account for other sparsity patterns of the outliers as well.
@@ 2.1 Algorithm and Main Result
Theorem 1.
$p\geq\frac{C_{\emph{sample}}}{\gamma^{2}}\cdot\frac{\kappa^{4}\mu^{3}r^{3}\log n}{n},\quad\alpha\leq\frac{c_{\emph{outlier}}}{\kappa^{2}\mu^{2}r^{2}}\cdot\frac{\gamma}{C_{\emph{thresh}}}$
and
for iteration $0\leq t\leq T$, where $T=n^{O(1)}$.
@@
Remark 3. Note that our result is also applicable for the noiseless case, i.e. the low rank matrix completion problem. In this case, we have improved the sample complexity of SVP from $O(\frac{\kappa^{6}\mu^{4}r^{6}\log n}{n})$ in [11] to $O(\frac{\kappa^{4}\mu^{3}r^{3}\log n}{n})$.
@@ 3 Proof Outline
$\max_{0\leq i\leq 2n}\left\|E^{t-1,i}\right\|_{2}:=\left\|E^{t-1,\infty}\right\|_{2}$ $\leq\frac{1}{C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\sigma_{r}^{\star}\right)\gamma^{t}$ (3a)
$\max_{0\leq i\leq 2n}\left\|\Delta^{t,i}\right\|_{2,\infty}:=\left\|\Delta^{t,\infty}\right\|_{2,\infty}$ $\leq\frac{5C_{10}}{C_{0}}\cdot\left(\frac{1}{\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t}$ (3b)
@@
$\leq$ $\frac{5C_{10}}{C_{0}}\left(\frac{1}{\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t}\cdot\left(2+\frac{5C_{10}}{C_{0}}\right)\sqrt{\frac{\mu r}{n}}\cdot\left(1+\frac{1}{C_{0}}\right)\sigma_{1}^{\star}+\frac{\mu r}{n}\cdot\frac{7}{C_{0}}\left(\frac{1}{\mu r}\sigma_{r}^{\star}\right)\gamma^{t}$
$\leq$ $\frac{12C_{10}}{C_{0}}\cdot\frac{1}{\mu r}\left(\frac{\mu r}{n}\sigma_{1}^{\star}\right)\gamma^{t}\leq\left(\frac{\mu r}{n}\sigma_{1}^{\star}\right)\gamma^{t},$
@@ 4 Initialization and Base Case
  • We use Lemma 4 instead of [11, Lemma 21] to bound $\left\|\mathcal{H}_{\Omega}\left(L^{\star}\right)\right\|_{2}$, and use Bernstein’s inequality applied to vectors instead of scalars to bound $\left\|e_{m}^{T}\mathcal{H}_{\Omega}^{(-i)}\left(L^{\star}\right)V^{1,m}\right\|_{2}$, yielding tighter bounds in terms of $\mu$ and $r$.
@@
Step 1: The Bound For $\left\|E^{0,\infty}\right\|_{2}$. First consider $\left\|E_{1}^{0,i}\right\|_{2}$. Recall from Assumption 3 that $\mathcal{P}_{\Omega}\left(S^{\star}\right)$ is $2\alpha p$-sparse. One has
$\left\|E_{1}^{0,i}\right\|_{2}=$ $\left\|p^{-1}\mathcal{P}_{\Omega}^{(-i)}\left(S^{0,i}-S^{\star}\right)\right\|_{2}\leq p^{-1}\cdot\left(2\alpha pn\right)\cdot C_{\text{thresh}}\frac{\mu r}{n}\sigma_{1}^{\star}\leq\frac{\gamma}{2C_{0}}\cdot\frac{1}{\kappa\mu r}\sigma_{r}^{\star},$ (4)
@@
$\alpha\leq\frac{1}{4C_{0}}\frac{1}{\kappa^{2}\mu^{2}r^{2}}\cdot\frac{\gamma}{C_{\text{thresh}}}.$
@@
$\leq$ $2c_{4}\sqrt{\frac{\mu r\log n}{np}}\sigma_{1}^{\star}\leq\frac{\gamma}{2C_{0}}\cdot\frac{1}{\kappa\mu r}\sigma_{r}^{\star},$
@@
$p\geq\frac{16c_{4}^{2}C_{0}^{2}}{\gamma^{2}}\cdot\frac{\kappa^{4}\mu^{3}r^{3}\log n}{n}.$
@@
$\left\|E^{0,i}\right\|_{2}\leq\left\|E_{1}^{0,i}\right\|_{2}+\left\|E_{2}^{0,i}\right\|_{2}\leq\frac{1}{C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\sigma_{r}^{\star}\right)\gamma.$ (6)
@@
$\left\|R\right\|_{2}\leq\left\|\Sigma^{\star}H^{1,i}-G^{1,i}\Sigma^{\star}\right\|_{2}\cdot\left\|\left(\Sigma^{\star}\right)^{-1}\right\|_{2}\leq\frac{4}{C_{0}}\cdot\left(\frac{1}{\mu r}\right)\gamma.$
@@
$\left\|T_{1}\right\|_{2}\leq\frac{4}{C_{0}}\cdot\left(\frac{1}{\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma.$
@@
$\left\|\left(\Sigma^{1,i}\right)^{-1}-\left(\Sigma^{\star}\right)^{-1}\right\|_{2}=\max_{1\leq k\leq r}\left|\frac{1}{\sigma_{k}^{1,i}}-\frac{1}{\sigma_{k}^{\star}}\right|=\max_{1\leq k\leq r}\frac{\left|\sigma_{k}^{1,i}-\sigma_{k}^{\star}\right|}{\sigma_{k}^{1,i}\sigma_{k}^{\star}}\leq\frac{2}{C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\right)\gamma\cdot\frac{1}{\sigma_{r}^{\star}}.$
@@
$\left\|T_{2}\right\|_{2}\leq\frac{2}{C_{0}}\cdot\left(\frac{1}{\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma.$
@@
$\left\|T_{3}\right\|_{2}\leq\frac{20}{C_{0}}\cdot\left(\frac{1}{\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma.$
@@
$\leq$ $\frac{1}{2C_{0}}\frac{\gamma}{\kappa\mu r}\sigma_{r}^{\star}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{1,\infty}\right\|_{2,\infty}\right),$
@@
$\leq$ $\frac{C_{10}}{2c_{4}C_{0}}\frac{\gamma}{\kappa\mu r}\sigma_{r}^{\star}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{1,\infty}\right\|_{2,\infty}\right),$
@@
$\left\|T_{4}\right\|_{2}\leq\frac{\gamma}{C_{0}}\left[\left(1+\frac{C_{10}}{c_{4}}\right)\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{1,\infty}\right\|_{2,\infty}\right)+\left\|D^{1,\infty}\right\|_{\mathrm{F}}\right].$
@@
$\left\|\Delta^{1,\infty}\right\|_{2,\infty}\leq\frac{17+\frac{C_{10}}{c_{4}}}{C_{0}}\left(\frac{1}{\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma+\frac{1+\frac{C_{10}}{c_{4}}}{C_{0}}\left\|\Delta^{1,\infty}\right\|_{2,\infty}+\frac{1}{C_{0}}\left\|D^{1,\infty}\right\|_{\mathrm{F}}.$ (8)
@@
$\leq$ $\frac{1}{2C_{0}}\frac{\gamma}{\kappa\mu r}\sigma_{r}^{\star}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{1,\infty}\right\|_{2,\infty}\right),$
@@
$\leq$ $\frac{1}{2C_{0}}\frac{\gamma}{\kappa\mu r}\sigma_{r}^{\star}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{1,\infty}\right\|_{2,\infty}\right),$
@@
$B_{2}=\left\|e_{m}^{T}\mathcal{H}_{\Omega}\left(L^{\star}\right)V^{1,m}\right\|_{2}\leq\frac{C_{10}}{2c_{4}C_{0}}\frac{\gamma}{\kappa\mu r}\sigma_{r}^{\star}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{1,\infty}\right\|_{2,\infty}\right),$
@@
$\leq$ $\frac{1}{2C_{0}}\frac{\gamma}{\kappa\mu r}\sigma_{r}^{\star}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{1,\infty}\right\|_{2,\infty}\right),$
@@
$\left\|\Delta^{1,\infty}\right\|_{2,\infty}$ $\leq\frac{17+\frac{C_{10}}{c_{4}}}{C_{0}}\left(\frac{1}{\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma+\frac{1+\frac{C_{10}}{c_{4}}}{C_{0}}\left\|\Delta^{1,\infty}\right\|_{2,\infty}$
@@
$\left(1-\frac{2+2C_{10}}{C_{0}}\right)\left\|\Delta^{1,\infty}\right\|_{2,\infty}\leq$ $\frac{18+2C_{10}}{C_{0}}\cdot\left(\frac{1}{\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma,$
$\left\|\Delta^{1,\infty}\right\|_{2,\infty}\leq$ $\frac{18+2C_{10}}{C_{0}-2-2C_{10}}\cdot\left(\frac{1}{\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma.$
@@
$\left\|\Delta^{1,\infty}\right\|_{2,\infty}\leq\frac{5C_{10}}{C_{0}}\cdot\left(\frac{1}{\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma.$